# Volumes, Lorenz-like Templates, and Braids

### v1

@@ Abstract.
Lorenz-like templates are obtained from applying half twists to the two strips of the Lorenz template. Ghrist proved that infinitely many of these templates are universal, meaning they contain all links in the 3-sphere. In this paper, we generalise the bunch algorithm to construct links embedded in universal Lorenz-like templates and provide an upper volume bound that is quadratic in the trip number. We also establish a correspondence between links embedded in these universal Lorenz-like templates and generalised T–links. While generalised T-links have appeared in the knot theory literature before, we prove that they are, in fact, all links in the 3-sphere. Thus, we can regard each link in the 3-sphere as a generalised T-link, which has an associated Lorenz link. We use the upper bound obtained from the generalised bunch algorithm to provide an upper volume bound for all hyperbolic link complements in the 3-sphere. Such a bound grows quadratically in terms of the braid indexes of the associated Lorenz links.
@@ 1. Introduction
We note that two generalised T-links with different set of parameters can represent the same link, see Theorem 4.8.
@@ 4. A new twist on Lorenz-like links
In this section, we provide a simpler braid description for all Lorenz-like links that can be embedded into the Lorenz-like templates $\mathscr{L}(0,v)$ and $\mathscr{L}(u,0)$, as obtained by Birman and Kofman for the Lorenz template in [2]. As a consequence, we obtain a simpler braid description for all links in the 3-sphere. This simpler braid description can be viewed as a generalisation of torus links and provides a simple parametrization for all links in the 3-sphere. Unexpectedly, Champanerkar, Futer, Kofman, Neumann, and Purcell have previously worked with this family of links in [6, Definition 1.2], but they didn’t realize that they are, in fact, all links in the 3-sphere, as we will demonstrate.
@@
If $n=0$, then $T^{0}$-links are the same as T-links, up to a number of unknots. This occurs because Birman and Kofman in [2] didn’t consider splittable links that have trivial components parallel to the left boundary component of the Lorenz template. This completely fine because the study of a splittable link is done by studying its components separately. However, we need to consider all components if we want to find all links in $\mathbb{S}^{3}$.
@@
Consider $\mathrm{Vol}(M)$ the sum of the volumes of the hyperbolic pieces of $M$. Since Dehn filling doesn’t increase volume by Thurston [21, Proposition 6.5.2], we have

### v2

@@ Abstract.
In this paper, we find a more straightforward problem that is equivalent to one of the major challenges in knot theory: the classification of links in the 3-sphere. More precisely, we provide a simpler braid description for all links in the 3-sphere in terms of generalised T-links. With this, we translate the problem of classifying links in the 3-sphere into a problem of counting the number of generalised T-links that represent the same link. Generalised T-links are a natural generalisation of twisted torus links and Lorenz links, two families of links that have been extensively studied by many people. Moreover, we generalise the bunch algorithm to construct links embedded in universal Lorenz-like templates and provide an upper volume bound that is quadratic in the trip number. We use the upper bound obtained from the generalised bunch algorithm for generalised T-links to establish an upper bound for the sum of the volumes of the hyperbolic pieces of all closed, orientable, connected 3-manifolds.
@@ 4. A new twist on Lorenz-like links
In this section, we provide a simpler braid description for all Lorenz-like links that can be embedded into the Lorenz-like templates $\mathscr{L}(0,v)$ and $\mathscr{L}(u,0)$, as obtained by Birman and Kofman for the Lorenz template in [2]. As a consequence, we obtain a simpler braid description for all links in the 3-sphere. This simpler braid description can be viewed as a generalisation of torus links and provides a simple parametrization for all links in the 3-sphere. Unexpectedly, Champanerkar, Futer, Kofman, Neumann, and Purcell have previously worked with this family of links in [6, Definition 1.2], but they did not realize that they are, in fact, all links in the 3-sphere, as we will demonstrate.
@@
If $n=0$, then $T^{0}$-links are the same as T-links, up to a number of unknots. This occurs because Birman and Kofman in [2] did not consider splittable links that have trivial components parallel to the left boundary component of the Lorenz template. This completely fine because the study of a splittable link is done by studying its components separately. However, we need to consider all components if we want to find all links in $\mathbb{S}^{3}$.
@@
Consider $\mathrm{Vol}(M)$ the sum of the volumes of the hyperbolic pieces of $M$. Since Dehn filling does not increase volume by Thurston [21, Proposition 6.5.2], we have